OpenStudy (anonymous):

the greater of two numbers is 8 more than 5 times the smaller. their sum is 68. Find the numbers?

OpenStudy (anonymous):

\[8+5x=y\] Where the "greater" is y. Thats how I read the first sentence. Think about how you'd read the second part.

OpenStudy (anonymous):

how do i know when I am finding y or x?

OpenStudy (anonymous):

y=8+5x x=y=68 x+8+5x=68 6x+8=68 6x=60

OpenStudy (anonymous):

x=10

OpenStudy (anonymous):

y=58

OpenStudy (anonymous):

Looks alright to me. I assume the line where x=y=68 is supposed to read x+y=68. You know what you are finding because its the value you don't have. Fortunately in this case you have two different relations for y and x... which means you can combine those relations to find them both.